# envname-P envname#1

### v2

@@ 4.2 Contract Metrics for Test Case
Let $A=\{a_{1},\ldots,a_{n}\}$ be the set of all contract assertions and $T=\{t_{1},\ldots,t_{m}\}$ be the generated test cases. $F_{t}\subseteq A$ is the set of violated assertions when executing a test case $t\in T$. Finally, $T_{neg}\subseteq T$ represents the set of test cases that successfully violated at least one contract.
Assert Violation Coverage (AVC)
AVC quantifies the coverage of assert statements that are successfully violated by all the test cases.
$AVC=\frac{|\{a_{i}\mid\exists t\in T_{neg}:a_{i}\in F_{t}\}|}{n}.$
The value of 1.0 for AVC ensures that all contract assertions are captured by the test cases. Lower values expose the unexplored regions of the contracts from the test cases.
Target Specificity (TS)
TS evaluates the precision of each test case by measuring how accurately it violates its intended set of target contracts. To formalize this, we first define $V\subseteq A$ as the set of contracts that a given test case $t\in T_{neg}$ is intended to violate. We then measure the alignment between this intended set $V_{t}$ and the actually violated set $F_{t}$ using the Jaccard Index. The final TS score is the average of these individual precision scores across all negative test cases:
$TS=\frac{1}{|T_{neg}|}\sum_{t\in T_{neg}}\frac{|F_{t}\cap V_{t}|}{|F_{t}\cup V_{t}|}.$
A score of 1.0 indicates perfect precision, where every test case violates exactly the set of contracts it was designed to. Lower scores reveal a discrepancy, indicating that test cases either failed to trigger their intended violations or caused unintended, collateral violations.
4.3 Contract-Aware Code Generation
This section details our methodology for systematically evaluating and enhancing the ability of LLMs to generate code snippets that robustly enforce contracts. To achieve a fine-grained understanding of a model’s contract awareness, we investigate its performance under two distinct prompting strategies. First, we establish a baseline condition, Contract Specification (CS). For this strategy, we use a powerful Commercial LLM to naturally integrate the contract rules from the HumanEval+ and MBPP+ datasets into the main text of the prompt. This creates a comprehensive prompt that describes both the functional goal and its contractual constraints in natural language. Second, we introduce an enhanced condition, which we term Example-Augmented Specification (EAS). This strategy builds upon the CS prompt by augmenting it with a single, precisely targeted contract- violating test cases for each described contract. This provides a concrete example of what constitutes a violation, intended to guide the model toward more robust code generation. A concrete example of this EAS prompt is shown in Figure 2.
To measure the impact of these prompting strategies, we assess the generated code snippets using a comprehensive suite of metrics for both functional correctness and contract adherence. Functional correctness is measured using the standard pass@k metric over a set of valid test cases. Contract adherence is evaluated from two perspectives: runtime enforcement using test cases and static alignment of the generated code snippets with the ground-truth contracts, as detailed in Section 4.4.
4.4 Contract Metric for Code Generation
For the evaluation of the generated code snippets, we employ pass@k to measure functional correctness on valid inputs and use AVC to measure the correctness of the generated contract assertions. We also design two additional metrics to more precisely evaluate contract-aware code generation.
Let $A=\{a_{1},\ldots,a_{n}\}$ be the set of ground-truth contract assertions, and let $\hat{A}=\{\hat{a}_{1},\ldots,\hat{a}_{m}\}$ be the set of assertions extracted from the LLM-generated code. Let $M\subseteq A\times\hat{A}$ be the set of pairs $(a_{i},\hat{a}_{j})$ where the ground-truth contract $a_{i}$ and the generated assertion $\hat{a}_{j}$ are determined to be semantically equivalent.
Assertion Alignment Recall (AAR)
AAR measures the model’s ability to implement all required contracts without omission. It is the proportion of ground-truth contracts that are successfully covered by at least one assertion in the generated code, functioning as a recall metric.
$AAR=\frac{|\{a_{i}\in A\mid\exists\hat{a}_{j}\in\hat{A}:(a_{i},\hat{a}_{j})\in M\}|}{n}.$
A high AAR score indicates that an output code ensures that all required contract specifications are generated.
Assertion Alignment Precision (AAP)
AAP measures the accuracy of the generated assertions, penalizing irrelevant or hallucinated ones. It is the proportion of generated assertions that correspond to a valid ground-truth contract, functioning as a precision metric.
$AAP=\frac{|\{\hat{a}_{j}\in\hat{A}\mid\exists a_{i}\in A:(a_{i},\hat{a}_{j})\in M\}|}{m}.$
A high AAP score indicates that the code does not contain assertions of unnecessary or incorrect checks.
@@ 5.1 Datasets
Our study utilizes HumanEval+ and MBPP+ benchmarks (Liu et al., 2023), which are enriched with assertion-level contracts. However, these benchmarks have a critical limitation for evaluating code robustness: the contract specifications are neither included in the model prompts, nor are contract-violating inputs included in the official test suites. This evaluation setup exclusively tests for functional correctness on well-formed inputs, leading to an inflated perception of code quality. We construct a supplementary dataset containing contract-violating test cases, as described in Section 4.1.
5.2 Candidate Models
We utilize o4-mini as our primary test case generator. For our code generation experiments, we evaluate a set of open-source models including gemma-3-12B-it (gemma-3), Deepseek-R1-Distill-Qwen-14B (DeepSeek-R1), Qwen3-14B (Qwen-3), and Phi-4-reasoning-plus (Phi-4-plus).
@@ 6 Empirical Studies
Our empirical study is structured over three research questions (RQs) designed to evaluate the PACT framework from multiple perspectives. The evaluation is conducted on HumanEval+ and MBPP+. We begin by assessing the quality and precision of the CVTs generated by PACT (RQ1). Next, we investigate whether providing these concrete test cases is more effective for eliciting contract-awareness than using abstract natural language descriptions alone (RQ2). Finally, we analyze the resulting trade-off between the enforced contract adherence and the functional correctness of the generated code snippets (RQ3).
Table 1: Evaluation results of the test case generation on HumanEval+ and MBPP+.
Benchmark Method AVC ($\uparrow$}) TS ($\uparrow$}) AVG ($\uparrow$})
HumanEval+ o4-mini 97.14% 75.60% 86.37%
o4-mini + SMT Solver 95.53% 85.81% 90.67%
MBPP+ o4-mini 94.67% 69.11% 81.89%
o4-mini + SMT Solver 93.50% 84.54% 89.02%
6.1 RQ1: How Effective is PACT in Generating High-Quality Contract- Violating Test Cases?
Our first research question investigates the effectiveness of our proposed framework, PACT, in generating high-quality CVTs. A naive baseline approach, which we refer to as o4-mini in Table 1, uses an LLM o4-mini for direct test case generation from given contracts. This method is often inadequate, as it disregards dependencies among contracts and produces logically inconsistent violations, making it difficult to precisely verify individual contract predicates.
In contrast, PACT (o4-mini+ SMT Solver) employs a more robust two-stage approach. First, an LLM converts contracts into rules for an SMT solver. Subsequently, an SMT solver uses these rules to generate test cases. This process guarantees the generation of logically sound test cases that are precisely targeted to violate a specific subset of contracts while adhering to the rest.
The empirical results in Table 1 validate the superiority of PACT. While both methods achieve high AVC scores, PACT significantly outperforms the direct generation in TS, achieving over 10%p increases on both HumanEval+ and MBPP+. This demonstrates that our two-stage approach successfully gives test cases precisely for intended contracts. The baseline’s slightly higher AVC is an expected outcome, as direct generation often triggers a wide range of assertion errors indiscriminately, artificially inflating coverage. PACT, on the other hand, focuses on valid, targeted violations.
While PACT generates valid rules, providing a precise test for the input contracts, there are still errors. A case-specific analysis of minor errors that occur even with well-formed rules is detailed in Appendix A.
Table 2: Evaluation results of contract adherence in code generation on HumanEval+.
Model Mode pass@1 ($\uparrow$}) AVC ($\uparrow$}) AAR ($\uparrow$}) AAP ($\uparrow$}) AVG ($\uparrow$})
gemma-3 CS 84.41% 24.85% 11.41% 14.04% 32.79%
EAS 82.94% 91.02% 28.07% 27.77% 57.45%
DeepSeek CS 73.78% 44.12% 15.65% 16.97% 37.63%
EAS 71.77% 79.29% 27.62% 28.01% 51.67%
Qwen3 CS 78.92% 28.04% 13.17% 22.55% 35.67%
EAS 77.83% 87.81% 31.53% 36.09% 58.31%
Phi-4-plus CS 72.23% 52.91% 18.78% 21.09% 41.25%
EAS 67.08% 69.50% 21.33% 20.06% 44.49%
Table 3: Evaluation results of contract adherence in code generation on MBPP+.
Model Mode pass@1 ($\uparrow$}) AVC ($\uparrow$}) AAR ($\uparrow$}) AAP ($\uparrow$}) AVG ($\uparrow$})
gemma-3 CS 78.56% 57.99% 17.50% 17.93% 41.49%
EAS 78.60% 95.57% 32.29% 31.82% 59.57%
DeepSeek CS 62.53% 64.20% 17.59% 17.57% 40.47%
EAS 60.15% 86.70% 28.23% 27.94% 45.47%
Qwen3 CS 72.41% 70.86% 21.09% 22.99% 46.84%
EAS 72.63% 94.85% 31.54% 32.30% 57.83%
Phi-4-plus CS 64.89% 67.33% 24.26% 24.65% 45.28%
EAS 63.76% 74.88% 29.20% 28.95% 49.20%
6.2 RQ2: Are Concrete Test Cases More Effective than Abstract Descriptions for Eliciting Contract-Awareness?
We compared two prompting methods across various models in Tables 2 and 3. CS, our baseline, provides only an abstract natural language description of the contracts. In contrast, EAS augments the CVTs, generated from Section 4.1, to the CS prompt.
The results demonstrate that for all of our base models, EAS is a more powerful method for eliciting contract-aware code generation. EAS achieves 33.7%p, 11.3%p, and 9.4%p improvements for AVC, AAR, and AAP, respectively, over CS. For models such as Qwen3, DeepSeek-R1, and Phi-4-plus, switching from CS to EAS leads to a dramatic improvement across all contract adherence metrics. This indicates that concrete examples of CVTs provide a clear and unambiguous signal that forces the model to move beyond purely functional logic. The increase in AAR and AAP indicates that the models are correctly generating test cases required for assertion checks, while the relatively huge improvements in AVC show that the resulting codes are more robust against a wider range of invalid inputs.
Concrete examples of contract violations provide a clear, unambiguous signal that forces the model to move beyond purely functional logic and implement the explicit enforcement of contracts—the predefined agreements on how ill-formed inputs must be rejected. For concrete examples of the code snippets and their corresponding LLM-generated test cases, please refer to Appendix B. However, this heightened focus on contract enforcement introduces a notable trade-off, as the model may generate more complex code that impacts its performance on purely functional correctness metrics. This trade-off is analyzed in further detail in the following section RQ3.
6.3 RQ3: How Does Enforcing Contract Adherence Impact the Functional Correctness of Generated Code?
While the RQ2 established that augmenting prompts with CVTs significantly enhances a model’s contract adherence, this section investigates the resulting trade-off with respect to its functional correctness. Our analysis reveals that a heightened focus on contract enforcement often comes at the cost of a measurable decline in performance on standard functional correctness benchmarks. This trade-off can be attributed to a shift in the model’s optimization focus. When provided with only a functional description, the LLM’s sole objective is to produce code snippets that pass functionality-focused tests. However, when the prompt is augmented with CVTs, the model must simultaneously satisfy two competing objectives: generating functionally correct logic and enforcing the specified contractual preconditions. This dual objective increases the complexity of the code generation task. The model is compelled to allocate part of its reasoning capacity to interpreting and implementing the contract rules, which can lead to subtle errors or oversights in the core functional logic. A concrete example of this is presented in Appendix C. Consequently, while the resulting code snippets are more robust and secure against invalid inputs, they may exhibit a lower pass@1 rate on test suites composed entirely of valid, well-formed inputs.
In conclusion, our findings reveal an inherent tension between contract adherence and functional correctness in LLM-based code generation. The pursuit of robustness through contract enforcement comes at a tangible cost to functionality. This trade-off highlights the critical need for advanced training paradigms—such as the reinforcement learning with multi-objective rewards proposed by PACT—capable of optimizing for both objectives simultaneously.
@@ 7 Conclusion
We introduce PACT, the first framework to redefine code and test case correctness by evaluating adherence to task specifications through both functionality and contract-based behavior. While prior benchmarks assess only pass@k on well-formed inputs, PACT introduces a comprehensive paradigm with dual test suites—one for functionality and one for contract violations—along with specific metrics to analyze contract awareness and uncover latent defects, enabling a more principled evaluation of code robustness. Our empirical evaluation demonstrates the effectiveness of PACT across multiple dimensions. We first show that PACT’s SMT-solver-based test case generation method ensures more accurate CVTs than direct generation with over 10%p better performance. Furthermore, our results reveal that augmenting prompts with these CVTs is a highly effective strategy for generating robust and contract-aware code, achieving 18.13%p increase in contract-specific metrics and 12.8%p in total average. Our empirical studies on various LLMs demonstrate the effectiveness of PACT, achieving approximately, but also uncovers a critical trade-off between this enhanced contract adherence and a model’s performance on functional correctness. These findings confirm that PACT provides a more complete and realistic assessment of an LLM’s contract-aware code generation capabilities, moving beyond the limitations of existing benchmarks.
8 Future work directions
While this work focuses on evaluating contract adherence, a natural next step is to actively improve it through advanced training methodologies. A promising direction is to leverage the novel metrics introduced in our PACT framework as direct training signals. The novel metrics introduced in our PACT framework, such as the runtime metric Assert Violation Coverage (AVC) and the static metrics Assertion Alignment Recall (AAR) and Precision (AAP), are particularly well-suited for this purpose and could be integrated into a multi-objective reward function for RL. This approach could enable models to learn to navigate the trade-off between functional correctness and contract adherence more effectively, optimizing for both objectives simultaneously.

### v3

@@ 4.3 ContractEval Metrics
Let $A=\{a_{1},\ldots,a_{n}\}$ be the set of contract assertions and $T=\{t_{1},\ldots,t_{m}\}$ be the generated test cases. For a test case $t\in T$, let $F_{t}\subseteq A$ denote the set of assertions violated when executing $t$. AVC measures the fraction of generated test cases that successfully trigger at least one assertion violation:
$\mathrm{AVC}=\frac{|\{t\mid t\in T\wedge|F_{t}|>0\}|}{|T|}.$
A higher AVC indicates that the generator more consistently produces CVTs that elicit at least one assertion failure. A lower AVC indicates that a larger portion of generated tests do not trigger any assertion failure.
TS evaluates how each CVT matches its intended contract targets. Let $T_{\mathrm{neg}}=\{t\in T\mid|F_{t}|>0\}$ be the set of tests that trigger at least one assertion violation. For each $t\in T_{\mathrm{neg}}$, let $V_{t}\subseteq A$ denote the set of assertions that $t$ is intended to violate. TS measures the Jaccard similarity between the intended target set $V_{t}$ and the actually violated set $F_{t}$ and averages it over $T_{\mathrm{neg}}$:
$\mathrm{TS}=\frac{1}{|T_{\mathrm{neg}}|}\sum_{t\in T_{\mathrm{neg}}}\frac{|F_{t}\cap V_{t}|}{|F_{t}\cup V_{t}|}.$
A TS of $1$ indicates that each test violates exactly its intended target set. A smaller TS indicates that a test misses some intended violations, triggers unintended violations, or both.
Contract Satisfaction Rate (CSR)
To evaluate contract enforcement against contract-violating inputs, we use verified CVTs obtained from our filtering process. Let $T_{\mathrm{ver}}=\{t_{1},\ldots,t_{k}\}$ denote the verified CVT set. For a generated code snippet $C$, let $T_{\mathrm{rej}}\subseteq T_{\mathrm{ver}}$ be the subset of inputs that $C$ rejects by raising an assertion failure or an explicit rejection. CSR quantifies the fraction of verified CVTs that are rejected by $C$:
$\mathrm{CSR}=\frac{|\{t\mid t\in T_{\mathrm{ver}}\wedge t\in T_{\mathrm{rej}}\}|}{|T_{\mathrm{ver}}|}.$
@@ 5.1 Datasets
We evaluate on ContractEval built on HumanEval+ and MBPP+ (Liu et al., 2023). For each task, we use the original well-formed tests to measure functional correctness and a set of verified CVTs, which we constructed, to measure contract satisfaction. We run all experiments on the filtered task set with specification-aligned contracts and feasible CVTs.
5.2 Prompting Strategies
We evaluate three prompt formats: standard, contract specification (CS), and example-augmented specification (EAS). The standard prompt uses the original natural-language query. CS uses the contract-aware rewrite from Section 4.2. EAS appends a small set of verified contract-violating tests (CVTs; Section 4.1) to CS as negative examples. We also evaluate Chain-of-Thought (CoT) (Wei et al., 2022) and a multi-turn prompting variant.
5.3 Evaluation Metrics
As described in Section 4.3 We evaluate contract-violating test generation using AVC and TS. For code generation, we report pass@k on well-formed tests and CSR on verified CVTs, and additionally use CodeBLEU Ren et al. (2020) and an LLM-as-judge score to assess the syntactic and semantic quality of the generated assertion checks.
5.4 Evaluated Models
We employ o4-mini as the primary engine for contract-violating test case generation, contract-aware query reconstruction, and as an LLM-as-judge. For our code generation experiments, we evaluate a set of open-source models including gemma-3-12B-it (gemma-3), Deepseek-R1-Distill-Qwen-14B (DeepSeek-R1), Qwen3-14B (Qwen-3), Phi-4-reasoning (Phi-4), and Phi-4-reasoning-plus (Phi-4-plus). We provide detailed implementation settings in Appendix A for completeness and reproducibility.
@@ 6 Empirical Studies
Our empirical study answers three research questions about ContractEval. We evaluate the validity of ContractEval by checking that reference assertions are aligned with the task specification and that CVTs precisely target selected contract assertions while remaining feasible under the others. Using ContractEval, we quantify a blind spot of standard code generation evaluation: models fail to reject CVTs. Finally, we study whether prompting strategies beyond our default setting can improve the trade-off between functional correctness and contract satisfaction.
6.1 Validating ContractEval: Contract Alignments and CVTs
ContractEval evaluates whether generated programs check assertion-level contracts—preconditions and validity constraints encoded as reference assertions—by testing them on CVTs (ill-formed inputs). This evaluation is valuable when (i) the reference assertions align with the task specification and (ii) the ill-formed tests precisely target intended contract clauses, rather than triggering arbitrary failures.
Table 1: Comparison of contract-violating test cases generated by o4-mini and by ContractEval (ours).
Method AVC ($\uparrow$}) TS ($\uparrow$}) Average ($\uparrow$})
o4-mini 97.58% 72.69% 85.13%
ContractEval (ours) 94.11% 84.92% 89.52%
While ContractEval brings assertions from HumanEval+ and MBPP+, they are not designed as a contract-evaluation benchmark and do not provide a guarantee that the assertions are semantically aligned with the task description. We perform a contract-quality check using o4-mini as an LLM judge that assesses whether the assertions reflect input constraints stated or implied by the task specification. Across 542 tasks that include reference assertions, 88.04% are judged specification-aligned. We filtered out programs with misaligned assertions for accurate downstream contract evaluation, as explained in Section 4.1. This filtering removes noise so that subsequent analyses reflect the model capability more precisely.
Table 2: Functional correctness and contract satisfaction in code generation. Adding contract descriptions (CS) helps, and adding a few ill-formed examples (EAS) yields large gains in contract satisfaction.
Model Instruction Functional Contract Satisfaction
pass@1 ($\uparrow$) CSR ($\uparrow$) CodeBLEU ($\uparrow$) LLM-as-judge ($\uparrow$)
DeepSeek-R1 Standard 80.81% 0.00% 0.00% 0.00%
CS 73.29% 27.24% 34.29% 53.05%
EAS 70.87% 52.66% 54.51% 81.49%
gemma-3 Standard 81.76% 0.00% 0.00% 0.00%
CS 74.75% 22.69% 29.97% 41.74%
EAS 69.56% 48.54% 48.88% 72.04%
Phi-4 Standard 75.39% 0.00% 0.00% 0.00%
CS 71.29% 40.71% 46.66% 69.33%
EAS 72.99% 50.73% 50.86% 75.63%
Phi-4-plus Standard 75.97% 0.00% 0.00% 0.00%
CS 70.83% 38.24% 44.38% 65.31%
EAS 73.21% 50.94% 51.04% 75.98%
Qwen-3 Standard 74.79% 0.00% 0.00% 0.00%
CS 70.62% 24.29% 35.05% 49.71%
EAS 69.65% 51.83% 53.30% 77.27%
We validate that ContractEval’s ill-formed tests accurately probe intended contracts. We compare o4-mini as a direct LLM baseline that generates ill-formed tests to our neuro-symbolic synthesis pipeline, ContractEval. Table 1 shows that both approaches achieve high AVC, but our approach yields substantially higher TS, indicating that ContractEval produces more precise and targeted CVTs for contract combinations.
While ContractEval shows high AVC performance, its score is not 100% accurate, indicating that additional filtering is mandatory for precise evaluation. A CVT should be ill-formed with respect to the contract while still being feasible as an execution input. Otherwise, a test may trigger unrelated errors that do not reflect contract satisfaction. Thus, we apply a feasibility filtering detailed in Section 4.1. We take intersecting tasks that pass both filtering, which results in 364 tasks remaining. These filtering processes remove tests that are ill-typed or semantically irrelevant and ensures that our evaluation reflects intended contract checking behavior. We provide a case analysis of detailed errors in Appendix C.
6.2 Contract-Aware Code Generation
Recall that we use ContractEval to evaluate how well LLMs generate contract-checking assertions that reject CVTs in addition to functional logic. Specifically, we measure functional correctness with pass@1 on well-formed tests and contract satisfaction on CVTs. We compare 1) a standard prompt with task description, 2) a CS prompt with both task and contract description, and 3) a EAS prompt of CS with a small set of CVT examples.
From Table 2, it is straightforward that a standard prompt for all base LLMs yields zero performance on contract evaluation. While the standard prompt achieves strong functional accuracy with 75–82% pass@1 scores, at the same time, it yields 0.00 scores on all contract satisfaction metrics. This shows that passing conventional well-formed tests does not imply that the generated program implements the intended input-validity checks. ContractEval makes this blind spot measurable: without explicit steering with contract information, models tend to implement functional logic while silently accepting invalid inputs, leading to missing or incomplete assertion logic.
We first examine the CSR performance of CS and EAS as CSR accurately evaluates whether the generated contracts successfully raises AssertionErrors for CVTs. Adding contract constraints in natural language descriptions, we achieve 22–41%p increases in CSR, indicating that models translate abstract constraints into assertion checks. However, EAS yields the largest gains in CSR by accomplishing 10–30%p over CS. On average, EAS achieves 50.94%p CSR improvement and maintains 92% of a pass@1 score, compared with the standard prompt.
While CSR is a precise and verifiable metric to check contract satisfaction of LLM-generated codes, we also provide CodeBLEU and LLM-as-judge scores for syntactic and semantic evaluation. Note that the standard prompt still yields 0.00 scores for these two metrics. EAS performs the best for these two metrics, gaining 13.65%p and 20.65%p performance improvements over CS.
Even among functionally correct generations, contract satisfaction remains far from perfect. For EAS, when we restrict evaluation to tasks where the generated solution achieves pass@1=1=1, the mean CSR stays in the 0.51–0.59 range across models (Appendix H). Thus, functional correctness and contract satisfaction are related but not equivalent objectives, reinforcing the need for a benchmark that evaluates both.
6.3 How do Prompt Strategies Affect Contract-Aware Code Generation?
After establishing that EAS is effective for eliciting contract-aware behavior, we analyze whether additional prompting techniques improve the trade-off between functional correctness and contract satisfaction. We consider (i) Chain-of-Thought (CoT) prompting and (ii) a refinement-style multi-turn strategy that first generates a solution with functional logic and then refines the code with contract information. Table 3 summarizes the impact of these techniques across our base LLMs.
While CoT slightly improves pass@1 by 1.2%p for gemma-3 and Qwen-3, it frequently degrades both pass@1 and CSR among other LLMs. More critically, CoT catastrophically fails on Phi-4 and Phi-4-plus. These models collapse to 0–4% on both metrics. This suggests that CoT is not appropriate in our setting, where the output must remain executable and simultaneously demonstrate the intended rejection behavior on CVTs.
In contrast, multi-turn is stable across models and improves functional correctness over EAS in most cases. In most cases, EAS achieves superior CSR performance to that of multi-turn, however, multi-turn achieves better pass@1 performance, improving 1.8%p over EAS on average. For instance, Phi-4-plus increases pass@1 from 73.21% to 77.42%, and DeepSeek-R1 increases from 70.87% to 72.96%. However, this gain is often accompanied by a modest drop in CSR by 2–7%p, indicating that refinement tends to prioritize preserving functional behavior over implementing every contract check implied by CVTs.
Table 3: Comparison of additional prompting strategies. Multi-turn recovers pass@1 while maintaining competitive contract satisfaction.
Model Prompting pass@1 ($\uparrow$}) CSR ($\uparrow$})
DeepSeek-R1 EAS 70.87% 52.66%
+ CoT 61.11% 35.00%
+ Multi-Turn 72.96% 45.44%
gemma-3 EAS 69.56% 48.54%
+ CoT 70.46% 50.37%
+ Multi-Turn 71.34% 49.26%
Phi-4 EAS 72.99% 50.73%
+ CoT 3.66% 1.91%
+ Multi-Turn 71.88% 46.58%
Phi-4-plus EAS 73.21% 50.94%
+ CoT 0.00% 0.00%
+ Multi-Turn 77.42% 48.52%
Qwen-3 EAS 69.65% 51.83%
+ CoT 71.13% 45.24%
+ Multi-Turn 71.70% 49.23%
Overall, EAS remains the strongest single-turn strategy for contract satisfaction, while a multi-turn strategy is the most reliable way to regain pass@1. Yet, none of the tested prompting variants achieves both high pass@1 and near-perfect CSR, highlighting and motivating future methods evaluated under ContractEval.
@@ 7 Conclusion
ContractEval reframes code generation evaluation by testing not only functional correctness on well-formed inputs, but also whether generated program reject CVTs via intended assertions. Built on HumanEval+ and MBPP+, ContractEval augments each task with contract-targeted CVTs. We propose a neuro-symbolic pipeline to construct precise CVTs where an LLM converts contract assertions into SMT-recognizable constraints and an SMT solver enumerates satisfiable violation combinations. ContractEval further introduces metrics to quantify (i) the precision of generated CVTs and (ii) contract satisfaction of generated programs. Empirically, we find that standard prompts for code generation yields 0% for contract satisfaction, failing to reject CVTs. This reveals a blind spot in conventional benchmarks. By contrast, EAS substantially improves contract satisfaction by 50.94% while preserving 92% of pass@1, compared to the standard prompt. Overall, ContractEval provides a benchmark and methodology for evaluating assertion-level contract checks as an essential axis of robustness in LLM-based code generation.
8 Limitations
ContractEval is designed to make assertion-level contract satisfaction measurable and comparable, but several practical constraints define its current scope and point to natural extensions.
Scope of contracts.
ContractEval derives contracts from reference assertions, which may be incomplete with respect to all validity constraints that developers intend. We mitigate misalignment by filtering tasks whose assertions are judged inconsistent with the natural-language specification, but ContractEval does not claim to exhaustively capture every admissible/invalid input boundary for each task.
Coverage of invalid inputs.
CVTs are clause-targeted and feasibility-checked, but they still represent a finite set of contract-violation patterns. They may not cover all semantically invalid inputs (e.g., distributional shifts, adversarial edge cases, or domain-specific constraints not expressed as assertions). Future versions could combine ContractEval’s targeted CVTs with complementary robustness tests (e.g., fuzzing-style perturbations) while preserving interpretability.
Cost and evaluation dependencies.
SMT enumeration and feasibility filtering introduce computational overhead compared to purely LLM-based test generation, and any auxiliary model-based judgments (e.g., contract–spec alignment checks or semantic scoring) may inherit bias from the judge model. Reducing reliance on model-based judgments and improving efficiency are practical directions for scaling ContractEval.
Appendix A Implementation Detail
[…]
Appendix B Hidden Contracts and Contract-Violating Tests
Natural-language programming tasks rarely specify only functional behavior. They also imply a boundary of admissible inputs—type restrictions, format requirements, range constraints, and other validity rules that developers take for granted. In this paper, we call these validity rules contracts and refer to contract-violating tests (CVTs) as inputs that violate one or more contract clauses (i.e., ill-formed inputs with respect to the intended input space).
Figure 4: Hidden contracts and contract-violating tests (CVTs). A typical LLM solution can pass prior benchmarks that only test well-formed inputs while silently accepting contract-violating inputs.
Figure 4 illustrates why CVTs matter for code generation evaluation. The task asks for has_dup(nums) that returns whether an input sequence contains duplicates. A typical LLM-generated solution correctly implements the core logic using a set, and it will pass standard unit tests that only cover well-formed inputs. However, the task description carries a hidden contract that nums is a list of numbers. Without explicit input checks, the same program may silently accept contract-violating inputs (e.g., a string or a mixed-type list) and still return a plausible output. This creates a blind spot: pass@k can certify a solution as “correct” without ever testing whether it rejects contract-violating inputs.
ContractEval makes this missing axis measurable by augmenting each task with CVTs derived from assertion-level contracts. Concretely, CVTs are designed so that a contract-checking implementation should raise an error (e.g., by triggering an assertion) instead of producing a normal output. Beyond evaluation, the same CVT examples can be reused as a lightweight steering signal: including a small number of CVTs in the prompt encourages models to generate explicit contract checks (assertions) alongside the functional logic.